rx2 has 0 coefficient in the first equationQUESTION 2 Consider the linear system 11 + 0.5X1 T1 12 0.5x2 + 13 0.25x3 X3 0.2 -1.425 2 = whose solution is (0.9, -0.8,0.7).
Answer #1

Answer Consider the given callabong x,-z = 0.2 0.5u +H ₂ -0.2503 = -1.425 x, -0.5U2tuz = 2 written in Cochcient mabil. 10.5 1I 0.2 Initial gauss (14,9, 2) = (0,0,0) st approximation - +(0.2-060) +)) = (0-2] (1425-0.5(0-2) +0-2500)) = 444.525) = (2-(0